# Empirical evaluation of normalizing flows in Markov Chain Monte Carlo

### v1

@@ 3.2.1 Autoregressive NFs
Coupling bijections receive an input $x$ and partition it into disjoint inputs $(x_{A},x_{B})$. The output $f_{i}(x)=y$ is partitioned into $(y_{A},y_{B})$ on the same dimensions. Part B stays constant ($y_{B}=x_{A}$) while $y_{A}$ is computed as $y_{A}=\tau(x_{A};\phi(x_{B}))$, where $\tau$ is a transformer – a bijection, parameterized with $\phi(x_{B})$. The function $\phi$ is a conditioner, which takes one of the vectors as input and predicts the parameters for $\tau$. Coupling bijections are autoregressive as each dimension of $y$ is a function of the corresponding preceding dimensions in $x$. $y_{A}$ trivially stays constant, while $y_{B}$ is transformed using the preceding $x_{A}$. For the same reason, their Jacobian is also triangular and its determinant can be computed efficiently. The function $\phi$ need not be bijective, so we make it a deep neural network, which makes $f_{i}$ expressive. The composition $f$ thus accurately models complex distributions provided $\phi$ is sufficiently complex and the number of bijections $f_{i}$ is sufficiently large (Draxler et al., 2024; Lee et al., 2021). We compose coupling bijections with permutations (also bijections) to avoid repeatedly using the same dimensions in parts A and B. This also means the dimension order can be arbitrary when analyzing the Jacobian.
MA bijections explicitly compute each output dimension as a function of all preceding input dimensions. This also holds for inverse autoregressive (IA) bijections, which are the inverses of MA bijections. This is in contrast to coupling bijections, which retain the autoregressive property by partitioning the input and thus cleverly ignoring specific dimensions. MA bijections achieve this with a Masked Autoencoder for Distribution Estimation (Germain et al., 2015, MADE) as the conditioner. MADE maps a $d$-dimensional input to a $d$-dimensional output with an autoencoder whose weights are masked to retain the autoregressive property between layers and thus in the entire neural network. All coupling bijection transformers are compatible with MA bijection transformers and vice-versa. The drawback of MA bijections is that their inverse pass requires $\mathcal{O}(d)$ operations. This results in poor scaling with data dimensionality when we need to perform both the forward and inverse passes.
@@
  • For transformers, we consider shift (Dinh et al., 2015, used in NICE), affine map (Dinh et al., 2017, used in Real NVP), linear rational spline (Dolatabadi et al., 2020, LRS), rational quadratic spline (Durkan et al., 2019, RQS), and invertible neural networks (Huang et al., 2018, used in NAF): NNdeep, NNdense, and NNboth, corresponding to neural networks that are (1) deep and thin, (2) shallow and dense, and (3) deep and dense.
@@ 3.2.2 Residual NFs
Residual bijections map an input according to $f_{i}(x)=x+g_{i}(x)$, where $g$ outputs a residual value. The transformation in residual NFs is a composition of such bijections. A sufficiently long composition ensures that $x$ can gradually be transformed into a desired data point using small residual values. We place the investigated residual NFs into one of two categories: architectures based on the matrix determinant lemma and contractive residual architectures that incrementally transform data with contractive maps.
For residual architectures based on the matrix determinant lemma, the determinant of $f_{i}$ can be efficiently computed via the following equation:
$\det\left(\frac{\partial f_{i}(x)}{\partial f}\right)=\det\left(I+W^{\top}A^{-1}V\right)\det A=\det\left(A+VW^{\top}\right).$
@@
We investigate three architectures following this lemma: planar flows (Rezende and Mohamed, 2015), Sylvester flows (Berg et al., 2018), and radial flows (Tabak and Turner, 2013; Rezende and Mohamed, 2015), with the following functions and determinants in each layer:
$g_{p}(x)$ $=v\sigma(w^{\top}x+b),$ $\det\left(\frac{\partial f_{p}(x)}{\partial x}\right)=I+s(x)vw^{\top};$
$g_{s}(x)$ $=V\sigma(W^{\top}x+b),$ $\det\left(\frac{\partial f_{s}(x)}{\partial x}\right)=I+S(x)VW^{\top};$
$g_{r}(x)$ $=\frac{\beta(x-x_{0})}{\alpha+r(x)},$ $\det\left(\frac{\partial f_{r}(x)}{\partial x}\right)=I+\frac{\beta}{\alpha+r(x)}\left(I-\frac{(x-x_{0})(x-x_{0})^{\top}}{r(x)(\alpha+r(x))}\right).$
The subscripts $p,s,r$ correspond to planar, Sylvester, and radial flows. We define $r(x)=||x-x_{0}||_{2},s(x)=\sigma^{\prime}(w^{\top}x+b),S(x)=\sigma^{\prime}(W^{\top}x+b)$, where $\sigma$ is the sigmoid function.
@@
Contractive residual architectures are also compositions of residual bijections. They use contractive maps as the functions $g_{i}$. A map is contractive with respect to a distance function $\delta$ if there exists a constant $L<1$ such that $\delta\left(f_{i}(x),f_{i}(y)\right)\leq L\delta(x,y)$ for any $x,y\in\mathbb{R}^{d}$. By the Banach fixed point theorem  (Behrmann et al., 2019), any contractive map has one fixed point $x_{*}=f_{i}(x_{*})$, which we obtain by starting with an arbitrary $x_{1}$ and repeatedly applying $x_{k+1}=f_{i}(x_{k})$. This lets us compute the inverse of $f_{i}$ as well. The update $x_{k+1}=x^{\prime}-g_{i}(x_{k})$ is guaranteed to converge to $x_{*}=f_{i}^{-1}(x^{\prime})$ for any starting point $x_{1}$. We can compute an unbiased estimate of the log determinant with the Hutchinson trace estimator (Hutchinson, 1989) within a power series (Behrmann et al., 2019):
$\log\left|\det\left(\frac{\partial f_{i}(z)}{\partial z}\right)\right|=\sum_{i=1}^{\infty}\frac{(-1)^{k+1}}{k}\mathrm{Tr}\left(\frac{\partial g_{i}(z)}{\partial z}\right)\approx\sum_{j=1}^{n}\frac{(-1)^{k+1}}{k}w_{j}^{\top}\frac{\partial g_{i}(z)}{\partial z}w_{j},$
@@
where $w_{j}\sim\mathcal{N}(0,I)$. We alternatively compute the series with the Russian roulette estimator (Chen et al., 2019):
@@
$\sum_{i=1}^{\infty}\frac{(-1)^{k+1}}{k}\mathrm{Tr}\left(\frac{\partial g_{i}(z)}{\partial z}\right)\approx\mathbb{E}_{n,w}\left[\sum_{k=1}^{n}\frac{(-1)^{k+1}}{k}w_{j}^{\top}\left(\frac{\partial g_{i}(z)}{\partial z}\right)^{k}w_{j}/P(N\geq k)\right].$
@@
We investigate the difference between the two by evaluating both i-ResNet that uses the power series estimator and the residual flow (Chen et al., 2019, ResFlow) that uses the Roulette estimator. Both construct $g_{i}$ as neural networks with spectral regularization to ensure $L<1$, which makes the maps contractive. We provide neural network parameterization details in Appendix B.4.
@@ 3.2.3 Continuous NFs
$z_{1}=z_{0}+\int_{t_{0}}^{t_{1}}g_{\phi}(t,z_{t}),\;\log\left|\det\left(\frac{\partial f^{-1}(z_{0})}{\partial z_{0}}\right)\right|=-\int_{t_{0}}^{t_{1}}\mathrm{Tr}\left(\frac{\partial g_{\phi}(t,z_{t})}{\partial z_{t}}\right),$ (5)
where the neural network $g_{\phi}:\mathbb{R^{d}}\rightarrow\mathbb{R^{d}}$ determines the ODE: $g_{\phi}(t,z_{t})=\frac{\partial z_{t}}{\partial t}$. To compute the latent point, we subtract the integral in Equation 5 (left) from both sides. Similarly, we omit the minus sign in Equation 5 (right) to obtain the log determinant of $f(z_{1})$ with respect to $z_{1}$. We compute the integrals using numerical solvers. The integrals for point and log determinant computation can be computed jointly by combining the two ODEs in $\mathbb{R}^{d}$ into a single one in $\mathbb{R}^{2d}$.
@@
  • CNFEuler, which approximates the integral with the Euler-Maruyama solver in 150 steps (Salman et al., 2018) and parameterizes $g_{\phi}$ as a time-independent feed-forward network.
  • CNFRK, which approximates the integral with the adaptive Runge-Kutta 4(5) solver (Grathwohl et al., 2018; Finlay et al., 2020) and parameterizes $g_{\phi}$ with a time-dependent neural network.
  • CNFRK(R), which is the same as CNFRK, but regularizes $g_{\phi}$ by the squared norm of the Jacobian of its transformation (Grathwohl et al., 2018).
@@
$\mathrm{Tr}\left(\frac{\partial f(x)}{\partial x}\right)\approx\frac{1}{n}\sum_{i=1}^{n}w_{i}^{\top}\mathrm{Tr}\left(\frac{\partial f(x)}{\partial x}\right)w_{i},w_{i}\sim\mathcal{N}(0,I),$
@@ 3.3 Benchmark target distributions
We consider synthetic and real-world target distributions in our benchmark, which we describe in the following sections. Our benchmark includes various commonly analyzed targets and partly overlaps with the recently described NF benchmark for SVI (Agrawal and Domke, 2024).
@@ 3.3.1 Synthetic targets
Hierarchical Bayesian models are common ways of modeling real-world phenomena and often use priors with spatially varying curvature (Grumitt et al., 2022, 2024). Sampling from such priors is complex, which leads to long MCMC runs when we have few data points for the likelihood. We include the 100D funnel and 100D Rosenbrock distributions in our benchmark to facilitate choosing sampling methods in this scenario. The funnel distribution and the unnormalized Rosenbrock log density are respectively defined as:
@@
$\log\widetilde{p}(x)$ $=-\sum_{d=1}^{D/2}s(x_{2d-1}^{2}-x_{2d})^{2}+(x_{2d-1}-1)^{2},$
where $D$ is even and $s>0$ is a fixed scale parameter.
@@
The unnormalized double well density is defined as $\log\widetilde{p}(x)=-\sum_{d=1}^{D}(x^{2}-4)^{2}$. We provide details on target definitions in Appendix D.3. We view these distributions as increasingly complex due to their growing number of components and modes. A successful sampling method will retain good performance regardless of the number of modes.
@@ 3.3.2 Real-world Bayesian model posteriors
We include diverse Bayesian model posterior distributions that describe real-world phenomena. These are also commonly used for MCMC benchmarking (Magnusson et al., 2024):
@@
  • 3003D stochastic volatility target, which models evaluation of derivative securities, such as options.
@@ 3.4 Evaluation methodology
We compare samplers and NFs in estimating the second moment of the target with a given computational budget. We use the squared bias of the second moment as the comparison metric, shortened as $b^{2}$. Using $b^{2}$ relates our experiments to other works in NFMC, as it is a commonly used metric in the field (see e.g., Hoffman et al., 2019; Grumitt et al., 2022).
When comparing methods across different targets, $b^{2}$ cannot be naively compared due to different true second moments. We thus opt for a rank standardization approach (Urbano et al., 2019). We rank different methods from minimum to maximum $b^{2}$ on each target. We compute standardized ranks (SR) on each target, then observe a method’s empirical average rank $\overline{r}$ and standard error of the mean $\hat{\sigma}$ as an uncertainty estimate. This ensures that all methods are comparable across all targets. We choose not to transform SR to the $[0,1]$ interval as in  (Urbano et al., 2019) because it would result in skewed uncertainties that are difficult to interpret. We provide a detailed definition of SR in Appendix E.
@@ 4 Results
In this section, we show our main sampler and NF comparison results. For every experiment, we warm up the sampler for 3 hours and sample for 8 hours to ensure each NF has enough time and data to attain a good fit. In Appendix A, we provide additional results regarding NF operation speed and differences between autoregressive NF components, as well as experiments with short NFMC runs. We provide all experiment configuration details in Appendix B.1.
@@ 4.1.1 Sampler comparison across all targets and NFs
In the gradient-free setting, IMH, Jump MH, and NeuTra MH all outrank MH. This suggests that preconditioning and jumps both improve MH dynamics. Finding a balance between jumps and MH transitions is preferable to independent jumps or standard MH, as indicated by Jump MH outperforming both IMH and MH. Both Jump HMC and Jump MH outrank their NeuTra counterparts, with Jump MH also being better than NeuTra HMC. Independent NF jumps enable better exploration than NF-based preconditioning.

### v2

@@ 3.2.1 Autoregressive NFs
Coupling bijections receive an input $x$ and partition it into disjoint inputs $(x_{A},x_{B})$. The output $f_{i}(x)=y$ is partitioned into $(y_{A},y_{B})$ on the same dimensions. Part B stays constant ($y_{B}=x_{A}$) while $y_{A}$ is computed as $y_{A}=\tau(x_{A};\phi(x_{B}))$, where $\tau$ is a transformer – a bijection, parameterized with $\phi(x_{B})$. The function $\phi$ is a conditioner, which takes one of the vectors as input and predicts the parameters for $\tau$. Coupling bijections are autoregressive as each dimension of $y$ is a function of the corresponding preceding dimensions in $x$. $y_{A}$ trivially stays constant, while $y_{B}$ is transformed using the preceding $x_{A}$. For the same reason, their Jacobian is also triangular and its determinant can be computed efficiently. The function $\phi$ need not be bijective, so we make it a deep neural network, which makes $f_{i}$ expressive. The composition $f$ thus accurately models complex distributions provided $\phi$ is sufficiently complex and the number of bijections $f_{i}$ is sufficiently large [20, 35]. We compose coupling bijections with permutations (also bijections) to avoid repeatedly using the same dimensions in parts A and B. This also means the dimension order can be arbitrary when analyzing the Jacobian.
MA bijections explicitly compute each output dimension as a function of all preceding input dimensions. This also holds for inverse autoregressive (IA) bijections, which are the inverses of MA bijections. This is in contrast to coupling bijections, which retain the autoregressive property by partitioning the input and thus cleverly ignoring specific dimensions. MA bijections achieve this with a Masked Autoencoder for Distribution Estimation [25, MADE] as the conditioner. MADE maps a $D$-dimensional input to a $D$-dimensional output with an autoencoder whose weights are masked to retain the autoregressive property between layers and thus in the entire neural network. All coupling bijection transformers are compatible with MA bijection transformers and vice versa. The drawback of MA bijections is that their inverse pass requires $\mathcal{O}(D)$ operations. This results in poor scaling with data dimensionality when we need to perform both the forward and inverse passes.
@@
  • For transformers, we consider shift [18, used in NICE], affine map [19, used in Real NVP ], linear rational spline [17, LRS], rational quadratic spline [15, RQS], and invertible neural networks [29, used in NAF]: NNdeep, NNdense, and NNboth, corresponding to neural networks that are (1) deep and thin, (2) shallow and dense, and (3) deep and dense.
@@ 3.2.2 Residual NFs
In residual architectures, $f$ is a composition of residual bijections $f_{i}$. These map an input according to $f_{i}(x)=x+g_{i}(x)$, where $g$ outputs a residual value. A sufficiently long composition ensures that $x$ can gradually be transformed into a desired data point using small residual values. We place the investigated residual NFs into two categories: architectures based on the matrix determinant lemma and contractive residual architectures that incrementally transform data with contractive maps.
The former contain bijections $f_{i}$, designed to have the Jacobian determinant equal to $\det\left(A+VW^{\top}\right)$, where $A\in\mathbb{R}^{D\times D}$ is invertible and $V,W\in\mathbb{R}^{D\times M}$. If computing $\det A$ and $A^{-1}$ is tractable and $M\ll D$, the determinant can be computed efficiently via the matrix determinant lemma:
$\det\left(\partial_{x}f_{i}(x)\right)=\det\left(A+VW^{\top}\right)=\det\left(I+W^{\top}A^{-1}V\right)\det A.$
@@
We investigate three architectures following this lemma: planar flows [43], Sylvester flows [9], and radial flows [48, 43]. Let $g_{i}^{(p)},g_{i}^{(s)},g_{i}^{(r)}$ denote the residual functions for Planar, Sylvester, and radial flows, respectively. Residual function definitions and the resulting Jacobian determinants for Planar, Sylvester, and radial flows are, respectively:
$g_{i}^{(p)}(x)$ $=v\sigma(w^{\top}x+b),$ $\det\left(\partial_{x}f_{i}^{(p)}(x)\right)=1+\sigma^{\prime}(w^{\top}x+b)w^{\top}v,$
$g_{i}^{(s)}(x)$ $=V\sigma(W^{\top}x+b),$ $\det\left(\partial_{x}f_{i}^{(s)}(x)\right)=\det\left(I+S(x)W^{\top}V\right),$
$g_{i}^{(r)}(x)$ $=\frac{\beta(x-x_{0})}{\alpha+r(x)},$ $\det\left(\partial_{x}f_{i}^{(r)}(x)\right)=\left(1+\frac{\alpha\beta}{(\alpha+r(x))^{2}}\right)\left(1+\frac{\beta}{\alpha+r(x)}\right)^{D-1},$
[…]
@@
Contractive residual architectures are compositions of residual bijections, which use contractive maps as functions $g_{i}$. A map is contractive with respect to a distance function $\delta$ if there exists a constant $L<1$ such that $\delta\left(f_{i}(x),f_{i}(y)\right)\leq L\delta(x,y)$ for any $x,y\in\mathbb{R}^{d}$. By the Banach fixed point theorem [8], any contractive map has one fixed point $x_{*}=f_{i}(x_{*})$, which we obtain by starting with an arbitrary $x_{1}$ and repeatedly applying $x_{k+1}=f_{i}(x_{k})$. This lets us compute the inverse of $f_{i}$ as well. The update $x_{k+1}=x^{\prime}-g_{i}(x_{k})$ is guaranteed to converge to $x_{*}=f_{i}^{-1}(x^{\prime})$ for any starting point $x_{1}$. We can compute an unbiased estimate of the log determinant with the Hutchinson trace estimator [31] within a power series [8]:
$\log\left|\det\left(\partial_{z}f_{i}(z)\right)\right|=\sum_{i=1}^{\infty}\frac{(-1)^{k+1}}{k}\mathrm{Tr}\left(\partial_{z}g_{i}(z)\right)\approx\sum_{j=1}^{n}\frac{(-1)^{k+1}}{k}w_{j}^{\top}\partial_{z}g_{i}(z)w_{j},$
@@
where $w_{j}\sim\mathcal{N}(0,I)$. We alternatively compute the series with the Russian roulette estimator [10]:
@@
$\sum_{i=1}^{\infty}\frac{(-1)^{k+1}}{k}\mathrm{Tr}\left(\partial_{z}g_{i}(z)\right)\approx\mathbb{E}_{n,w}\left[\sum_{k=1}^{n}\frac{(-1)^{k+1}}{k}w_{j}^{\top}\left(\partial_{z}g_{i}(z)\right)^{k}w_{j}/P(N\geq k)\right],$
@@
where $N$ is a positive random variable and $n\sim N$. We use $N\sim\mathrm{Geom}(0.5)$ as in the original paper. We note that the Hutchinson trace estimator is also valid for $w_{j}^{(k)}\sim_{\mathrm{iid}}\mathrm{Rademacher}$, however, an analysis by [10] found the differences to be fairly small and even in favor of Gaussian random variables under a certain parametrization. We investigate the difference between the two by evaluating both i-ResNet that uses the power series estimator and the residual flow [10, ResFlow] that uses the Roulette estimator. Both construct $g_{i}$ as neural networks with spectral regularization to ensure $L<1$, which makes the maps contractive. We provide neural network parameterization details in Appendix B.4.
@@ 3.2.3 Continuous NFs
$z_{1}=z_{0}+\int_{t_{0}}^{t_{1}}g_{\phi}(t,z_{t}),\;\log\left|\det\left(\partial_{z_{0}}f^{-1}(z_{0})\right)\right|=-\int_{t_{0}}^{t_{1}}\mathrm{Tr}\left(\partial_{z_{t}}g_{\phi}(t,z_{t})\right),$ (5)
where the neural network $g_{\phi}:\mathbb{R^{d}}\rightarrow\mathbb{R^{d}}$ determines the ODE: $g_{\phi}(t,z_{t})=\partial_{t}z_{t}$. To compute the latent point, we subtract the integral in Equation 5 (left) from both sides. Similarly, we omit the minus sign in Equation 5 (right) to obtain the log determinant of $f(z_{1})$ with respect to $z_{1}$. We compute the integrals using numerical solvers. The integrals for point and log determinant computation can be computed jointly by combining the two ODEs in $\mathbb{R}^{d}$ into a single one in $\mathbb{R}^{2d}$.
@@
  • CNFEuler, which approximates the integral with the Euler-Maruyama solver in 150 steps [47] and parameterizes $g_{\phi}$ as a time-independent feed-forward network.
  • CNFRK, which approximates the integral with the adaptive Runge-Kutta 4(5) solver [22, 21] and parameterizes $g_{\phi}$ with a time-dependent neural network.
  • CNFRK(R), which is the same as CNFRK, but regularizes $g_{\phi}$ by the squared norm of the Jacobian of its transformation [22].
@@
$\mathrm{Tr}\left(\partial_{x}f(x)\right)\approx\frac{1}{n}\sum_{i=1}^{n}w_{i}^{\top}\partial_{x}f(x)w_{i},w_{i}\sim\mathcal{N}(0,I),$
@@ 3.3 Benchmark target distributions
We consider synthetic and real-world target distributions in our benchmark, which we describe in the following sections. Our benchmark includes various commonly analyzed targets and partly overlaps with the recently described NF benchmark for SVI [2].
@@ 3.3.1 Synthetic targets
Hierarchical Bayesian models are common ways of modeling real-world phenomena and often use priors with spatially varying curvature [24, 26]. Sampling from such priors is complex, which leads to long MCMC runs when we have few data points for the likelihood. We include the 100D funnel and 100D Rosenbrock distributions in our benchmark to facilitate choosing sampling methods in this scenario. The funnel distribution and the Rosenbrock log density are respectively defined as:
@@
$\log p_{X}(x)$ $=-\sum_{d=1}^{D/2}s(x_{2d-1}^{2}-x_{2d})^{2}+(x_{2d-1}-1)^{2}-C,$
where $D$ is even, $s>0$ is a fixed scale parameter, and $C$ is the log of the normalization constant.
@@
The double well density is defined as $\log p_{X}(x)=-\sum_{d=1}^{D}(x^{2}-4)^{2}-C$, where $C$ is the log of the normalization constant. We provide details on target definitions in Appendix D.3. We view these distributions as increasingly complex due to their growing number of components and modes. A successful sampling method will retain good performance regardless of the number of modes.
@@ 3.3.2 Real-world Bayesian model posteriors
We include diverse Bayesian model posterior distributions that describe real-world phenomena. These are also commonly used for MCMC benchmarking [39]:
@@
  • 3003D stochastic volatility target, which models the evaluation of derivative securities, such as options.
@@ 3.4 Evaluation methodology
We compare samplers and NFs in estimating the second moment of the target with a given computational budget. We use the squared bias of the second moment as the comparison metric, shortened as $b^{2}$. It measures the difference between the true second moment and the second moment as estimated using MCMC samples. We provide a detailed definition and discuss its relation to the bias-variance decomposition of mean squared error in Appendix E. Using $b^{2}$ relates our experiments to other works in NFMC, as it is a commonly used metric in the field [see e.g., 30, 24].
When comparing methods across different targets, $b^{2}$ cannot be naively compared due to different true second moments. We thus opt for a rank standardization approach [49]. We rank different methods from minimum to maximum $b^{2}$ on each target. We compute standardized ranks (SR) on each target, then observe a method’s empirical average rank $\overline{r}$ and standard error of the mean $\hat{\sigma}$ as an uncertainty estimate. This ensures that all methods are comparable across all targets. We choose not to transform SR to the $[0,1]$ interval as in [49] because it would result in skewed uncertainties that are difficult to interpret. We provide a detailed definition of SR in Appendix E.
@@ 4 Results
In this section, we show our main sampler and NF comparison results. For every experiment, we warm up the sampler for 3 hours and sample for 8 hours to ensure each NF has enough time and data to attain a good fit. In Appendix A, we provide additional results regarding NF operation speed and differences between autoregressive NF components, as well as experiments with short NFMC runs. We provide all experiment configuration details in Appendix B.1. We also repeat some analyses using kernelized Stein discrepancy [34] as the comparison metric in Appendix A.4, which describes other properties of empirical MCMC sample distributions. In Appendix A.5, we also evaluate Jump MCMC with an iterated sampling importance resampling kernel for global proposals [46], which uses multiple proposals that further improve Jump MCMC efficiency.
@@ 4.1.1 Sampler comparison across all targets and NFs
Figure 1: Numerical comparison of NFMC methods on the entire benchmark.
@@
In the gradient-free setting, IMH, Jump MH, and NeuTra MH all outrank MH. This suggests that preconditioning and jumps both improve MH dynamics. Finding a balance between jumps and MH transitions is preferable to independent jumps or standard MH, as indicated by Jump MH outperforming both IMH and MH. A pure global proposal strategy appears to outperform classic MCMC when the underlying transition kernel dynamics are not very expressive. This can explain IMH ranking better than MH, but worse than HMC. Indeed, [7] similarly found IMH to yield a better effective sample size and Kolmogorov-Smirnov statistic values than the less-expressive MALA sampler on synthetic and real-world examples.
Both Jump HMC and Jump MH outrank their NeuTra counterparts, with Jump MH also being better than NeuTra HMC. These results thus suggest that, on average, independent NF jumps enable better exploration than NF-based preconditioning. This is consistent with a field theory analysis [23], where a purely global NF sampler outperforms NeuTra MCMC with the local MALA and elliptical slice sampler kernels. In some other experiments, the authors investigated NeuTra MCMC, Jump MCMC, and the combination thereof, and found examples where each performs the best. Our experiments on this benchmark clarify that the greatest benefits stem from independent NF jumps.
@@
We note that SR compares the investigated samplers relative to each other and is not an absolute measure of sample quality. For example, HMC achieves a worse SR in Figure 1(b) than in Figure 1(a) because of differently parametrized NFMC samplers. Its $b^{2}$ is the same in both plots, but the NFMC $b^{2}$ values are different, which results in different SR values.